Uxx + Uyy =0 convert the situation equation into its Canonical form and find out its general solution


1
Expert's answer
2021-11-23T08:37:12-0500

this is the canonical form of elliptic equation, 2-D Laplace equation:

"u_{xx} + u_{yy}=0"


we have "u_{xx} =- u_{yy}=a"

where a is constant


general solution:

"u(x, y) = \frac{a}{4}(x^2-y^2)"


solution for rectangle with sides L and H:


for boundary conditions:

"u(0,y)=g(y),u(L,y)=0,u(x,0)=0,u(x,H)=0"


"u(x,y)=\sum B_n sinh(\frac{n\pi (x-L)}{H})sin(\frac{n\pi y}{H})"


"B_n=\frac{2}{Hsinh(\frac{n\pi (-L)}{H})}\int^H_0 g(y)sin(\frac{n\pi y}{H})"


for boundary conditions:

"u(0,y)=0,u(L,y)=0,u(x,0)=0,u(x,H)=f(x)"


"u(x,y)=\sum B_n sinh(\frac{n\pi y}{L})sin(\frac{n\pi x}{L})"


"B_n=\frac{2}{Lsinh(\frac{n\pi H}{L})}\int^L_0 f(x)sin(\frac{n\pi x}{L})"